# A Super Version of a theorem of Fricke-Klein

### v1

@@ 3.2. Supergroups
Let $G$ be superscheme and consider $m:G\times G\rightarrow G$ to be a morphism of schemes. Furthermore, if there exists morphisms of superschemes
@@
then we call $G$ a group superscheme. If G is a $k$-superscheme, we call it an algebraic supergroup.
@@
We notice, that since an algebraic supergroup is defined by limit diagrams (in this case pullbacks), that whenever we have a limit preserving functor (e.g. the inclusion or bosonic reduction as both of them are right adjoints), that group objects will be sent to group objects.
Another standard way to think about superschemes, is in terms of representable functor
@@
Let $\mathcal{C}$ be a locally small category and let $\operatorname{Set}$ be the category of sets. For each object $A$ of $\mathcal{C}$ let $\operatorname{Hom}(A,-)$ be the hom functor that maps object X to the set $\operatorname{Hom}(A,X)$.
A functor
(10) $\mathcal{F}:\mathcal{C}\rightarrow\operatorname{Set}$
is said to be representable if it is naturally isomorphic to Hom(A,–) for some object A of C. A representation of F is a pair $(A,\phi)$ where
(11) $\phi:\operatorname{Hom}(-,A)\rightarrow\mathcal{F}$
is a natural isomorphism.
@@
Affine schemes are representable by a functor from $\operatorname{Sch}^{\operatorname{op}}\rightarrow\operatorname{Set}$. Furthermore, in this setting one can think of an algebraic group as a functor from
@@
(13) $v=\sum\alpha_{i}\otimes e_{i},$
so we were able to write down a vector, with odd coefficients, which can not be seen, by considering the pedestrian definition given earlier. This is important, because to prove a later result we have to work with $R$-points of matrices and vectors.
@@
because we ring morphisms are determined by the evaluation maps for each variable. The evaluations live inside of $\mathbb{C}$ and have to fulfill the determinant condition. The result is precisely $\operatorname{SL}_{2}(\mathbb{C})$.
However, for now let us return to the first definition of a group superscheme. Assume $G$ to be an affine $k$-superscheme and apply the spec functor to this series of diagrams and we get the following commuting diagrams:
@@
If we now live in the affine setting then it also turns out that we have a functor from
$\operatorname{SAlg}_{k}\rightarrow Grp.$
An easy consequence of this remark is, that for a $k$ Grassman algebra $\Lambda$ we have the natural quotient map from $\Lambda\rightarrow k$. We use the fact, that we indeed have a group functor which results in the induced morphism
$G(\Lambda)\rightarrow G(k)=G_{red}(k).$
Another important notion in the theory of algebraic supergroups, while not immediately relevant here, is the notion of a Harish-Chandra pair. We choose to write this short section as we believe it is quite helpful in general when thinking about algebraic supergroups. It turns out that this category is equivalent to the category algebraic supergroups [15] [9].
@@
Let SHCP be the category of super Harish-Chandra pairs and ASP the category of algebraic supergroups then as mentioned earlier we also have the following equivalence of categories
(14) $\textbf{SHCP}\cong\textbf{ASP}$
@@
The supergroups whose group structure is preserved under the splitting, were also already classified. It turns out it is equivalent to $\mathfrak{g}_{1},\mathfrak{g}_{1}]=0 [$24].
@@
The group law is given by ordinary matrix multiplication. Alternatively, we can also describe it as a group preserving a super symplectic structure $\omega$ on $\mathbb{R}^{2|1}$ [18], so there can arise confusion due to this $\operatorname{OSp}(1|2)$ acting $\mathbb{R}^{2|1}$. The descriptions are equivalent and for now we will proceed with the symplectic one. Generally we will be working over $\mathbb{C}$ and denote this by $\operatorname{OSp(1|2)}$. We can also think of $\operatorname{OSp}(1|2)$ as a group functor from the category of algebraic superschemes to the category of groups. We recover the previous notion of $\operatorname{OSp}(1|2)$ matrices via the functor of points and thinking of the $\Lambda$-points as matrices of the form
@@
with $a,b,c,d,f\in\Lambda_{0}$ and $\alpha,\beta,\gamma,\delta\in\Lambda_{1}$. These matrices are also subject to some relations, but before we write them out, we should also know where they come from. Since we are defining a group preserving the symplectic form, we can also write out the symplectic form as a matrix which is
$J=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix}$
which should be invariant under the action of $\operatorname{OSp}(1|2)$. From this we also have a nice formula for the inverse of an element $g$. Let $g\in\operatorname{OSp}(1|2)$, then its inverse is given by
@@
$g=\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}\quad,\quad g^{-1}=\begin{pmatrix}d&-b&\delta\\ -c&a&-\gamma\\ -\beta&\alpha&f\end{pmatrix}$
@@
onto us and we impose the relation $\operatorname{Ber}(g)=1$ separately. One can also notice, that the underlying bosonic group is $\operatorname{SL}(2,-)$. It follows that we naturally get a surjective ring homomorphism induced by quotienting all odd generators and we get a closed immersion of schemes
@@
i.e.
(18) $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\longmapsto\begin{pmatrix}a&b&0\\ c&d&0\\ 0&0&1\end{pmatrix}$
By the functorial nature of algebraic groups, we can now also consider the natural morphism from a $k$-Grassmann algebra $\Lambda$ to $k$, by sending all odd generators to 0. This induces the morphism
$\operatorname{OSp}_{1|2}(\Lambda)$$\operatorname{OSp}_{1|2}(k)=SL_{2}(k)$}
and the identification can be done with the functor of points in the same way as we did earlier for the functor $SL_{2}(-)$, so in formulas we have
(19) $\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}\longmapsto\begin{pmatrix}a_{\#}&b_{\#}&0\\ c_{\#}&d_{\#}&0\\ 0&0&1\end{pmatrix}\longmapsto\begin{pmatrix}a_{\#}&b_{\#}\\ c_{\#}&d_{\#}\end{pmatrix}$
Furthermore, when we take a look at the composition
@@
We get a splitting for the supergroup $\operatorname{OSp}_{1|2}$. We can once again introduce notions like semisimple or reductive, however being a semisimple algebraic supergroup is a very special phenomena. The classification of semisimple algebraic supergroups is well known and was originally proven by Hochschild and Djokovic [7]. Recently Sherman reproved this with more modern techniques in [21]. We will state Shermans version.
@@
Next we describe the mentioned triangulation lemma for $\operatorname{OSp}(1|2)$.
@@
Let $A_{0}\in SL(2,\mathbb{C})\subset\operatorname{OSp}(1|2)$ and $B_{0}\in\operatorname{OSp}(1|2)$. Furthermore, assume that $A$ can be diagonalized as $A=diag(\mu,\mu^{-1},1)$ and $\mu\neq\mu^{-1}$, then one can always find an element $g\in\operatorname{OSp}(1|2)$ that acts by simultaneous conjugation on the pair $(A,B)$ such that
@@
On the other hand we now have to conjugate the matrix $B_{0}$:
@@
where we remind ourselves that $x=y^{-1}(\lambda-\lambda^{-1})^{-1}$ and $y$ is a solution of the polynomial $P$. We are working in a non-integral ring, so the existence of a solution is non-obvious. However, a result of Shigeaki Nagamichi [12], tells us as long as the reduction has roots, the corresponding super polynomial also has roots. ∎
@@
In Witten and Stanford [23] they make use of this parametrization of the moduli space of flat $\operatorname{OSp}(1|2)$-connections to compute an explicit form of $G$-invariant measures. They also write it down in the form
@@
$q_{1}$ $=\begin{pmatrix}0&0&1\\ 0&0&0\\ 0&-1&0\end{pmatrix}$ $q_{2}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 1&0&0\end{pmatrix}$
This might look different compared to the parametrization given earlier, however one remembers that the Taylor expansion of odd things vanishes after the first order. Expanding it that way gives you the desired result. Furthermore, one notes that this parametrization of our pair $(A,B)$ ensures that there are 3 even degrees of freedom and 2 odd degrees of freedom (defined up to sign). We also note, that $q_{1},\ q_{2}$ are the generators of $\mathfrak{g}_{1}$ of $\mathfrak{osp}(1|2)$ which you simply exponentiate to turn them into elements in the corresponding algebraic supergroup.
@@ 3.4. Invariant Theory
In this section, we want to introduce the necessary tools from classical invariant theory and adapt them to the $\mathbb{Z}_{2}$-graded case. Most of these results will generalize without adapting the proof. We follow [13] and [17]
First a short remark on $\operatorname{OSp}(1|2)$. Since we have semisimple representations for $\operatorname{OSp}(1|2)$ we also get that it is linearly reductive. We start by showing the so-called Reynolds operator exists, which is important for the following decomposition.
@@
Let us employ the outline strategy in a specific example, i.e. let us reprove the result of Fricke-Klein again.
@@
One should be careful though, that this isomorphism is an isomorpism of modules and not of rings. However, it does give us enough information about the nature and number of the generators which is enough for us.
@@
It follows that $v\in W^{\prime}$, so it is non-empty invariant and $W^{\prime}\cap V^{G}=0$. By applying Zorn’s lemma to the set of invariant subspaces, with this property we can find a maximal one, which we are gonna call $V^{\prime}$. Suppose, there are two such maximal subspaces, which we are going to denote by $V_{1},V_{2}$. It holds for the direct sum that $V_{1}\oplus V_{2}\cap V^{G}=0$. To show that, assume that it is non-empty, thus we can find an element $w=v_{1}\oplus v_{2}$, which is invariant under the group action. However, since the action acts on both parts separately, this is a contradiction, hence $V_{1}\oplus V_{2}\cap V^{G}=0$. This means, that the direct sum, would be a larger subspace with the wanted property, which once again is a contradiction. Thus $V^{\prime}$ is uniquely determined and we can also see that
@@
Now we define $R$ by the property, that it fixed $V^{G}$ and has $V^{\prime}$ as its kernel. ∎
@@
Consider a direct sum $V=V_{1}\oplus...\oplus V_{r}$ of finite dimensional super vector spaces. A function $f\in k[V_{1}\oplus...\oplus V_{r}]$ is called mukltihomogeneous of degree $h=(h_{1},...,h_{r})$ if $f$ is homogeneous of degree $h_{i}$ in $V_{i}$, i.e. for all $v_{1},...,v_{r}\in V$ and $t_{1},...,t_{r}\in k$ we have
@@
Every polynomial function $f$ is in a unique way a sum of multihomogeneous functions $f=\sum f_{h}$. The $f_{h}$ are usually called multihomogenous components of $f$. This gives rise to a decomposition
@@
It is clear that the multidegree is preserved under the action of $\operatorname{GL}(V_{1})\times...\times\operatorname{GL}(V_{r})$. Furthermore, if $V_{1},...V_{r}$ are representations, than we also have a natural grading on the ring of invariants
@@
where $f_{s_{1},...,s_{d}}\in k[V^{d}]$ are well-defined and multihomogeneous of multidegree $(s_{1},...,s_{d})$. We call $f_{1,...,1}\in k[V^{d}]$ the (full) polarization of $f$ denoted by $\mathcal{P}f$
@@
It remains to show, that the pairings are still given by traces, which is essentially just a retelling of Procesi [20], but with the word super attached.
@@
This follows by a simple calculation. Take a homogeneous basis $\left\{e_{j}\right\}$ for $V$ and let $\left\{\varphi_{j}\right\}$ be its dual basis. We take a vector $v\otimes\varphi\in V\otimes V^{\vee}$ and compute the coefficients of the associated matrix $A$. Let
@@
Let $V$ be a super vector space. Then the space multilinear functions on $V^{m}\oplus(V^{\vee})^{m}$ is isomorphic to the space of multilinear functions of $\operatorname{End}(V)^{m}$
@@
We immediately recognize, that the right hand side are the multilinear functions on $\operatorname{End}(V)^{m}$. ∎
@@
This was also proven by Lehrer and Zhang in [14] with a different method of proof earlier than the stated result mostly in the language of Brauer diagrams. With that method they also computed the polynomial invariants ones from the multilinear ones, which are simply the monomials in the multilinear invariants.
@@
4. Character Varieties
We follow [22]. Let $G$ be a reductive algebraic group over k. When looking at the set $Hom(\Gamma,G)$, we can associate to it a ring $\mathcal{R}(\Gamma,G)$, and the $k$-points of $\operatorname{|}{Spec}(\mathcal{R}(\Gamma,G))|=\operatorname{Hom}(\Gamma,G)$. Since every finitely generated group can be constructed by imposing a set of relation on the free group $F_{n}$, we will start by constructing the representation variety of $F_{n}$.

### v2

@@ 3.2. Algebraic Supergroups
Let $G$ be an $S$-superscheme and consider $m:G\times G\rightarrow G$ to be a morphism of schemes. Furthermore, if there exists morphisms of superschemes
@@
then we call $G$ an $S$-group superscheme, where we take the fiber products over $S$. If G is a $k$-superscheme, we call it an algebraic supergroup.
@@
We notice that since an algebraic supergroup is defined by limit diagrams (in this case pullbacks), that whenever we have a limit preserving functor, that group objects will be sent to group objects.
Let $\mathcal{C}$ be an arbitrary category with all small limits, then we call an object $G\in\mathcal{C}$ that satisfies all these conditions a group object. Since the category of superschemes is complete, we obtain, by cheating, an equivalent definition.
@@
An algebraic supergroup $G$ is a group object in the category of superschemes.
@@
Generally, one denotes the group objects inside of a category $\mathcal{C}$ that contains all small limits by $\operatorname{Grp}(\mathcal{C})$.
Another standard way to think about superschemes, is in terms of representable functor, in particular, in this setting one can think of an algebraic group as a functor from
@@
(14) $v=\sum\alpha_{i}\otimes e_{i}.$
@@
because the morphisms are determined by the evaluation maps for each variable. The evaluations live inside of $\mathbb{C}$ and have to satisfy the determinant condition. The result is precisely $\operatorname{SL}_{2}(\mathbb{C})$.
However, for now let us return to the first definition of a group superscheme. Assume $G$ to be an affine $k$-superscheme and apply $\mathrm{Spec}(-)$ to this collection of diagrams, and we obtain the following:
@@
By the functor of points viewpoint a superscheme is already determined by its functor on affines, i.e.
$\operatorname{sRings}\rightarrow\mathrm{Set}.$
If $G$ is an algebraic group over $k$ and $\Lambda$ a $k$-Graßman algebra, then by functoriality we obtain a map
$G(\Lambda)\rightarrow G(k)=G_{red}(k),$
which arises out of the natural quotient map of $\Lambda\rightarrow k$ that sends all odd elements to zero.
Another important notion in the theory of algebraic supergroups, while not immediately relevant here, is the notion of a Harish-Chandra pair. We choose to write this short section as we believe it is quite helpful in general when thinking about algebraic supergroups. It turns out that this category is equivalent to the category algebraic supergroups [11], [20].
@@
Let $\mathrm{SHCP}$ be the category of super Harish-Chandra pairs and $\mathrm{ASP}$ the category of algebraic supergroups then as mentioned earlier we also have the following equivalence of categories
(15) $\mathrm{SHCP}\cong\mathrm{ASP}$
@@
The supergroups whose group structure is preserved under the associated graded functor, were also already classified. It turns out it is equivalent to $\mathfrak{g}_{1},\mathfrak{g}_{1}]=0 [$32].
@@
The group law is given by ordinary matrix multiplication. Alternatively, we can also describe it as a group preserving a supersymplectic structure $\omega$ on $\mathbb{R}^{2|1}$ [24]. The descriptions are equivalent and for now we will proceed with the symplectic description. Generally, we will be working over $\mathbb{C}$ and denote this by $\operatorname{OSp(1|2)}$. We can also think of $\operatorname{OSp}(1|2)$ as a group functor from the category of algebraic superschemes to the category of groups. We recover the previous notion of $\operatorname{OSp}(1|2)$-matrices via the $\Lambda$-points of $\mathrm{OSp}(1|2)$, where $\Lambda$ is a $\mathbb{C}$-Graßmann algebra. Explicitly, the $\Lambda$-points are given by matrices of the following form:
@@
with $a,b,c,d,f\in\Lambda_{0}$ and $\alpha,\beta,\gamma,\delta\in\Lambda_{1}$.
Remark 3.27.
This viewpoint is implicitly taken in [25].
These matrices are also subject to some relations, which can be extracted from the symplectic viewpoint for $\mathrm{OSp}(1|2)$. The symplectic form that has to be preserved can be written as a matrix $J$ of the form:
$J=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix}$,
which needs to be invariant under the action of $\operatorname{OSp}(1|2)$. We also obtain a formula for the inverse of an element $g$. Let $g\in\operatorname{OSp}(1|2)$, then its inverse is given by
@@
$g=\begin{pmatrix}a&b&\alpha\\ c&d&\beta\\ \gamma&\delta&f\end{pmatrix}\quad\quad g^{-1}=\begin{pmatrix}d&-b&\delta\\ -c&a&-\gamma\\ -\beta&\alpha&f\end{pmatrix}$.
@@
onto us and the relation $\operatorname{Ber}(g)=1$ is imposed separately.
Remark 3.28.
We recall that the functor $(-)_{\mathrm{red}}$ preserves group objects, and one observes that the underlying bosonic group is $\operatorname{SL}_{2}(-)$. The reduction functor induces a morphism of schemes, which is a closed immersion, i.e.
@@
Let $\iota:k\rightarrow\Lambda$ be the inclusion morphism and $f:\Lambda\rightarrow k$. By functoriality, we obtain the following commutative diagram:
$\mathrm{SL}_{2}(\Lambda)$$\mathrm{OSp}_{1|2}(\Lambda)$$\mathrm{OSp}_{1|2}(k)=\mathrm{SL}_{2}(k)$$\mathrm{SL}_{2}(k)$$\mathrm{OSp_{1|2}(k)}$$\scriptstyle{f_{*}}$$\scriptstyle{\iota_{*}}$$\scriptstyle{\iota_{*}}$
and observe the splitting sequence
@@
which in formulas is given by
(20) $\begin{pmatrix}a_{\#}&b_{\#}\\ c_{\#}&d_{\#}\end{pmatrix}\longmapsto\begin{pmatrix}a_{\#}&b_{\#}&0\\ c_{\#}&d_{\#}&0\\ 0&0&1\end{pmatrix}\longmapsto\begin{pmatrix}a_{\#}&b_{\#}\\ c_{\#}&d_{\#}\end{pmatrix}$
We can once again introduce notions like semisimple or reductive, however being a semisimple algebraic supergroup is a very special phenomenom. The classification of semisimple algebraic supergroups is well known and was originally proven by Hochschild and Djoković [9]. Recently Sherman reproved this with more modern techniques in [29]. We will state Sherman’s version.
@@
Next we describe the mentioned triangulation lemma for $\operatorname{OSp}(1|2)$. Let $\Lambda_{i}$ be the free odd superalgebra in $i$-generators. We will work in $\Lambda$-points, where $\Lambda\coloneqq\underset{i}{\mathrm{colim}}\ \Lambda_{i}$.
@@
Let $A_{0}\in SL(2,\mathbb{C})\subset\operatorname{OSp}(1|2)$ and $B_{0}\in\operatorname{OSp}(1|2)$. Furthermore, assume that $A$ can be diagonalized as $A=\mathrm{diag}(\mu,\mu^{-1},1)$ and $\mu\neq\mu^{-1}$, then one can always find an element $g\in\operatorname{OSp}(1|2)$ that acts by simultaneous conjugation on the pair $(A,B)$ such that
@@
On the other hand, we have to conjugate the matrix $B_{0}$:
@@
where we remind ourselves that $x=y^{-1}(\lambda-\lambda^{-1})^{-1}$ and $y$ is a solution of the polynomial $P$. We are working in a non-integral ring, so the existence of a solution is non-obvious. However, a result of Shigeaki Nagamichi [14], tells us as long as the reduction possesses roots, the corresponding super polynomial also has roots. ∎
@@
In [31] they make use of this parametrization of the moduli space of flat $\operatorname{OSp}(1|2)$-connections to compute an explicit form of $G$-invariant measures. They also write down the pair in the form
@@
$q_{1}$ $=\begin{pmatrix}0&0&1\\ 0&0&0\\ 0&-1&0\end{pmatrix}$ $q_{2}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 1&0&0\end{pmatrix}$.
This might look different compared to the parametrization given earlier, but one recalls that the Taylor expansion of odd elements vanishes after the first order. Expanding it that way gives you the desired result. Furthermore, one notes that this parametrization of our pair $(A,B)$ ensures that there are 3 even degrees of freedom and 2 odd degrees of freedom (defined up to sign). We also note that $q_{1},\ q_{2}$ are the generators of $\mathfrak{g}_{1}$ of $\mathfrak{osp}(1|2)$ and by exponentiation they become elements in the corresponding algebraic supergroup.
@@ 3.4. Invariant Theory
In this section, we begin by reproving the Fricke-Klein case again with a method that directly aligns with the method of our main result. This explicit worked example should help elucidate the proof strategy for the supercase. Afterwards, we introduce the necessary tools from classical invariant theory and adapt them to the $\mathbb{Z}_{2}$-graded case. Most of these results will generalize without adapting the proof and can be skipped for the experienced reader. We follow [15] and [23].
First a short remark on $\operatorname{OSp}(1|2)$. Since we have semisimple representations for $\operatorname{OSp}(1|2)$ we also get that it is linearly reductive. We start by introducing the notion of a Reynolds operator.
@@
Let us employ the outline strategy in a specific example, i.e. let us reprove the theorem of Fricke-Klein again.
@@
One should be careful though, that this isomorphism is an isomorphism of modules and not of rings. However, it does give us enough information about the number of generators which suffices.
@@
It follows that $v\in W^{\prime}$, so it is non-empty invariant and $W^{\prime}\cap V^{G}=0$. By applying Zorn’s lemma to the set of invariant subspaces, with this property we can find a maximal one, which we denote by $V^{\prime}$. Suppose that there are two such maximal subspaces denoted by $V_{1},V_{2}$. It holds for the direct sum that $V_{1}\oplus V_{2}\cap V^{G}=0$. To show that, assume that it is non-empty, thus we can find an element $w=v_{1}\oplus v_{2}$, which is invariant under the group action. However, since the action acts on both parts separately, this is a contradiction, hence $V_{1}\oplus V_{2}\cap V^{G}=0$. This means, that the direct sum, would be a larger subspace with the wanted property, which once again is a contradiction. Thus $V^{\prime}$ is uniquely determined and we can also see that
@@
Now we define $R$ by the property, that it fixes $V^{G}$ and has $V^{\prime}$ as its kernel. ∎
@@
Remark 3.39.
One can prove, that the existence of a Reynolds operator is equivalent to the representation category being semisimple. It is also equivalent to the fixed point functor/invariants functor $(-)^{G}$ being exact.
@@
Consider a direct sum $V=V_{1}\oplus...\oplus V_{r}$ of finite dimensional super vector spaces. A function $f\in k[V_{1}\oplus...\oplus V_{r}]$ is called mltihomogeneous of degree $h=(h_{1},...,h_{r})$ if $f$ is homogeneous of degree $h_{i}$ in $V_{i}$, i.e. for all $v_{1},...,v_{r}\in V$ and $t_{1},...,t_{r}\in k$ we have
@@
Every polynomial function $f$ is in a unique way a sum of multihomogeneous functions $f=\sum f_{h}$. The $f_{h}$ are usually called multihomogeneous components of $f$. This gives rise to a decomposition
@@
It is clear that the multidegree is preserved under the action of $\operatorname{GL}(V_{1})\times...\times\operatorname{GL}(V_{r})$. Furthermore, if $V_{1},...V_{r}$ are representations, then we also have a natural grading on the ring of invariants
@@
where $f_{s_{1},...,s_{d}}\in k[V^{d}]$ are well-defined and multihomogeneous of multidegree $(s_{1},...,s_{d})$. We call $f_{1,...,1}\in k[V^{d}]$ the (full) polarization of $f$ denoted by $\mathcal{P}f$.
@@
It remains to show, that the pairings are still given by traces, just like in [26].
@@
We prove this by calculation. Take a homogeneous basis $\left\{e_{j}\right\}$ for $V$ and let $\left\{\varphi_{j}\right\}$ be its dual basis. We take a vector $v\otimes\varphi\in V\otimes V^{\vee}$ and compute the coefficients of the associated matrix $A$. Let
@@
Let $V$ be a super vector space. Then the space of multilinear functions on $V^{m}\oplus(V^{\vee})^{m}$ is isomorphic to the space of multilinear functions on $\operatorname{End}(V)^{m}$
@@
We immediately recognize that the right hand side are the multilinear functions on $\operatorname{End}(V)^{m}$. ∎
@@
This was also proven by Lehrer and Zhang in [16] with a different method of proof earlier than the stated result mostly in the language of Brauer diagrams. With that method they also computed the polynomial invariants ones from the multilinear ones, which given by the monomials in the multilinear invariants.
@@
4. Main Result
We can now state our main result. Let $V$ be a $1|2$-dimensional super vector space.
Theorem 4.1.
Proof.
Let us consider the algebra
(53) $\mathbb{C}[\operatorname{End}(V)^{2}]^{\operatorname{OSp}(V)}.$
By (3.52) we know that the number of polynomial invariants of $V^{\otimes 4}$ is 10. By (3.55) we know that there is one relation for $\operatorname{dim}(V)=1|2$ and thus 9 independent generators. Because the multilinear invariants of matrices and vectors are the same, the polynomial invariants are also the same. By reductivity we have
[…]
$\alpha=b\gamma-a\delta,\quad\beta=d\gamma-c\delta,$
$\gamma=a\beta-c\alpha,\quad\delta=b\beta-d\alpha,$
$f=1+\alpha\beta,\quad f^{-1}=ad-bc,$
together with the condition $\operatorname{Ber}(A)=1$. We know that there exist no odd invariants on $\mathbb{C}[\operatorname{End}(V)^{2}]$ and we also know the co-action on $I$ sends the odd generators to odd generators. Thus all the odd relations, do not give us an invariant of $I$. However, the relation $\operatorname{Ber}(A)=1$ is invariant under the co-action, thus it is an invariant vector in $I$. It follows that
We denote by $\mathrm{gen}(-)$ the number of generators of a ring. Then
(56) $\begin{split}\mathrm{gen}(\mathbb{C}[\operatorname{End}(V)^{2}]^{\operatorname{OSp}(V)})&=\mathrm{gen}(\mathbb{C}[\operatorname{OSp}(V)^{2}]^{\operatorname{OSp}(V)})+\mathrm{gen}(I^{\operatorname{OSp}(V)})\\ 9&=\mathrm{gen}(\mathbb{C}[\operatorname{OSp}(V)^{2}]^{\operatorname{OSp}(V)})+2\\ \mathrm{gen}(\mathbb{C}[\operatorname{OSp}(V)^{2}]^{\operatorname{OSp}(V)})&=7\end{split}$
By an (3.48), we know that all multilinear invariants are given by traces, thus all these invariants are traces as well. ∎
5. Geometric Interpretations
5.1. Character Varieties
We follow [30]. Let $G$ be a reductive algebraic group over k. When looking at the set $\mathrm{Hom}(\Gamma,G)$, we can associate to it a ring $\mathcal{R}(\Gamma,G)$, and the $k$-points of $\operatorname{Spec}(\mathcal{R}(\Gamma,G))|=\operatorname{Hom}(\Gamma,G)$. Since every finitely generated group can be constructed by imposing a set of relations on the free group $F_{n}$, we will start by constructing the representation variety of $F_{n}$.